OpenStudy (anonymous):

|–4(-2) – 8| + |–1 – (-2^2)| + 2(-2)^2

OpenStudy (ybarrap):

$$ |–4(-2) – 8| + |–1 – (-2^2)| + 2(-2)^2\\ =|8 – 8| + |–1 – 4| + 2(4)\\ =|0| + |-5| + 8\\ =5+8\\ =13 $$

OpenStudy (ybarrap):

Let me know if you have any questions.

OpenStudy (anonymous):

thank you so much!

OpenStudy (ybarrap):

yw